OpenStudy (anonymous):

2x-[2x - {2x - (2x -2) - 2} - 2] - 2 HEEEEEEEEEELLLLLLLLLLLPPPPPPPPPPPP

OpenStudy (ankit042):

start by solving inner most bracket

OpenStudy (anonymous):

plz help me to get the simplification.. thanks so much

OpenStudy (imtiaz7):

2x-[2x - {2x - (2x -2) - 2} - 2] - 2 first we solve the small bracket =2x-[2x-{2x-2x+2-2}-2]-2 =2x-[2x-(0)-2]-2 =2x-[2x-2]-2 =2x-2x+2-2 =0

OpenStudy (anonymous):

woow... your so great ... thanks so much @Imtiaz7

OpenStudy (imtiaz7):

no problem.

OpenStudy (anonymous):

\[\sqrt[2](x4 y2)\] help me also here to simplify

OpenStudy (imtiaz7):

bherly first tell me is is x^4 or x4

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

x^4 y^2

OpenStudy (imtiaz7):